OpenStudy (anonymous):

3m - 2n + 4p = 15 m - n + p = 3 m + 4n - 5p = 0 System of equations

OpenStudy (anonymous):

Hello there, Discord. Rainbow Dash is here to help. Isolate any variable in the first equation then plug it to either second or third equation. Does this make sense?

OpenStudy (anonymous):

But I made you lose your loyalty and took your wings.

OpenStudy (anonymous):

And I could use some clarification

OpenStudy (anonymous):

Saying, I want to isolate m in first equation.\[3m - 2n + 4p = 15\]\[3m = 15 + 2n - 4p\]\[m = \frac{15 + 2n - 4p}{3} = 5 + \frac{2}{3}n - \frac{4}{3}p\] Ok, now I know m. Plug it into second or third equation. Does that make sense?

OpenStudy (anonymous):

Wait what does m equal?

OpenStudy (anonymous):

\[m = 5 + \frac{2}{3}n - \frac{4}{3}p\]

OpenStudy (anonymous):

Oh okay.

OpenStudy (anonymous):

You know what to do?

OpenStudy (anonymous):

I think so. Lemme try and solve it.